the points (2,3) and (6,11) liw on a line.
what is the growth rate of this line?

Answer by fractalier(6550)

If by growth rate, you mean slope (which is what I think you mean), then we can find it by
slope m = (change in y)/(change in x)
slope m = (11 - 3)/(6 - 2) = 8/4 = 2